OpenStudy (anonymous):

*Proportions, Ratios and Rates* Justin invented a zapper that shrinks and enlarges things. After shrinking almost everything in the house to 40% of its original size, Justin's mother demanded that he return things to normal. What percent settings should he use to zap objects back to their original size? (Show Work)

OpenStudy (unklerhaukus):

\[1\times\frac{40}{100}\times\frac?{100}=1\]